If we are constructing a 90% confidence interval for the variance of a population using a...

  1. If we are constructing a 90% confidence interval for the variance of a population using a sample of size 16, the two critical values for chi-square from the table that we need to use are  and   (use the three decimal place accuracy that the table gives you)

wondering if I have to take 16-1 for this question.
Answer #1

Solution :

Given that,

n = 16

df = 16 - 1 = 15

Two critical values are

2L = 2/2,df = 24.996

2R = 21 - /2,df = 7.261